# Decoding Time Series with LLMs: A Multi-Agent Framework for Cross-Domain Annotation

### v1

@@ I.2 Qualitative Examples
Table 15: A case study in Finance dataset. The target multivariate time series data is shown in Figure 5(d) (a). GPT-4o is the LLM backbone. Purple highlights basic patterns, green highlights general patterns, yellow highlights domain-specific patterns and blue highlights correlations between variables in multivariate time series data.
@@
Figure 5: Case study: a multivariate time series data from the Finance dataset, which has four variables, i.e., price, volumn, RSI and SMA.
@@
Table 19: One more example of general annotation generation in the Environment dataset. The selected dataset is shown in Fig. 7 (a). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.
@@
Table 20: One more example of general annotation generation in the Environment dataset. The selected dataset is shown in Fig. 7 (b). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.
@@
Table 21: One more example of general annotation generation in the Energy dataset. The selected dataset is shown in Fig. 8 (a). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.
@@
Table 22: One more example of general annotation generation in the Energy dataset. The selected dataset is shown in Fig. 8 (b). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.
@@
Table 23: One more example of general annotation generation in the Social Good dataset. The selected dataset is shown in Fig. 9 (a). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.
@@
Table 24: One more example of general annotation generation in the Social Good dataset. The selected dataset is shown in Fig. 9 (b). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.

### v2

@@ I.2 Additional Results on Other LLM Backbones
Table 23: A case study in Stock dataset. The target multivariate time series data is shown in Fig. 8(d) (a). GPT-4o is the LLM backbone. Purple highlights basic patterns, green highlights general patterns, yellow highlights domain-specific patterns and blue highlights correlations between variables in multivariate time series data.
@@
Figure 8: Case study: a multivariate time series data from the Stock dataset, which has four variables, i.e., price, volumn, RSI and SMA.
@@
Table 27: One more example of domain-specific annotation generation in the Environment dataset. The selected dataset is shown in Fig. 10 (a). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.
@@
Table 28: One more example of domain-specific annotation generation in the Environment dataset. The selected dataset is shown in Fig. 10 (b). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.
@@
Table 29: One more example of domain-specific annotation generation in the Energy dataset. The selected dataset is shown in Fig. 11 (a). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.
@@
Table 30: One more example of domain-specific annotation generation in the Energy dataset. The selected dataset is shown in Fig. 11 (b). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.
@@
Table 31: One more example of domain-specific annotation generation in the Social Good dataset. The selected dataset is shown in Fig. 12 (a). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.
@@
Table 32: One more example of domain-specific annotation generation in the Social Good dataset. The selected dataset is shown in Fig. 12 (b). Purple highlights basic patterns, green highlights general patterns, and yellow highlights the domain-specific patterns.